yeetantor:

3x + 1 over 4y2 What is the value of the expression above when x = 3 and y = 4? You must show all work and calculations to receive full credit.

snowflake0531:

\[\frac{ 3x+1 }{ 4y^2 }\] where \[x=3 \]\[y=4\]

snowflake0531:

So we have \[\frac{ 3(3)+1 }{ 4(4^2) }\] can you simplify that

Florisalreadytaken:

slow down -- first of all is it \[ \frac{3x + 1} {4y^2} \] or \[ 3x + \frac {1} {4y^2} \]

Florisalreadytaken:

you there mate? @yeetantor

snowflake0531:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Florisalreadytaken slow down -- first of all is it \[ \frac{3x + 1} {4y^2} \] or \[ 3x + \frac {1} {4y^2} \] \(\color{#0cbb34}{\text{End of Quote}}\) those two don't equal....

Florisalreadytaken:

exactly

Florisalreadytaken:

because it its the second one, we just failed to answer the thing properly lol

snowflake0531:

It's obviously the first one-

Florisalreadytaken:

if*

Florisalreadytaken:

it didnt say "ALL OVER"